### v1

# Finite-time Scaling beyond the Kibble-Zurek Prerequisite: Driven Critical Dynamics in Strongly Interacting Dirac Systems

Abstract
In conventional quantum critical point (QCP) characterized by order parameter fluctuations, the celebrated Kibble-Zurek mechanism (KZM) and finite-time scaling (FTS) theory provide universal descriptions of the driven critical dynamics. However, in strongly correlated fermionic systems where gapless fermions are usually present in vicinity of QCP, the driven dynamics has rarely been explored. In this Letter, we investigate the driven critical dynamics in two-dimensional Dirac systems, which harbor semimetal and Mott insulator phases separated by the QCP triggered by the interplay between fluctuations of gapless Dirac fermions and order-parameter bosons. By studying the evolution of physical quantities for different driving rates through large-scale quantum Monte Carlo simulation, we confirm that the driven dynamics is described by the FTS form. Accordingly, our results significantly generalize the KZM theory by relaxing its requirement for a gapped initial state to the system accommodating gapless Dirac fermionic excitation. Through successfully extending the KZM and FTS theory to Dirac QCP, our work not only brings new fundamental perspective into the nonequilibrium critical dynamics, but also provides a novel theoretical approach to fathom quantum critical properties in fermionic systems.
Introduction— Fathoming nonequilibrium universal properties near a quantum critical point (QCP) is one of the central issues in modern physics Dziarmaga (2010); Polkovnikov et al. (2011); D’Alessio et al. (2016); Mitra (2018). Although the general organizing principle for the nonequilibrum critical dynamics is still elusive, a unified framework for understanding the generation of topological defects after the linear quench was proposed by Kibble in cosmological physics and then generalized by Zurek in condensed matter systems  Kibble (1976); Zurek (1985). This celebrated Kibble-Zurek mechanism (KZM) has aroused intensive investigations from both theoretical and experimental aspects, exerting far-reaching significance in both classical and quantum phase transitions Kibble (1976); Zurek (1985); Laguna and Zurek (1997); Hindmarsh and Rajantie (2000); Chuang et al. (1991); Dziarmaga (1998); Zurek et al. (2005); Dziarmaga (2005); Damski and Zurek (2007); Lamporesi et al. (2013a); Navon et al. (2015); Du et al. (2023); Ko et al. (2019); Ulčakar et al. (2020); Crowley et al. (2020); Gómez-Ruiz et al. (2020); Goo et al. (2021); Maegochi et al. (2022); Dziarmaga et al. (2022); Sherman et al. (2023); Balducci et al. (2023); Zeng et al. (2023); Lamporesi et al. (2013b); Barends et al. (2016); King et al. (2022); Keesling et al. (2019); Ebadi et al. (2021); Yao et al. (2022); Qiu et al. (2020); Ebadi et al. (2022); Sunami et al. (2023); Li et al. (2023). More interestingly, it was found that scaling behaviors can also manifest themselves in the whole driven process Zhong and Xu (2005); Deng et al. (2009); De Grandi et al. (2010); Chandran et al. (2012); Clark et al. (2016); Kolodrubetz et al. (2012). As a generalization of the KZM, a finite-time scaling (FTS) theory was proposed to systematically understand the full scaling properties Gong et al. (2010); Feng et al. (2016). These full scaling forms have been verified in various systems from numerical to experimental works Huang et al. (2014); Liu et al. (2014); Yin et al. (2014); Liu et al. (2015); Yin et al. (2017); Li et al. (2019); Shu et al. (2023); Clark et al. (2016); Gong et al. (2010); Feng et al. (2016); King et al. (2023); Keesling et al. (2019); Ebadi et al. (2021); Garcia and Chepiga (2024); King et al. (2024). Moreover, the KZM and the FTS recently show their fabulous power in state preparations and probing critical properties in fast-developing programmable quantum devices King et al. (2023); Keesling et al. (2019); Ebadi et al. (2021); Garcia and Chepiga (2024); King et al. (2024); Dupont and Moore (2022).
The fermionic QCP Sachdev (2011); Hertz (1976); Millis (1993); Berg et al. (2019); Löhneysen et al. (2007); Lee (2018a); Metlitski and Sachdev (2010); Yerzhakov and Maciejko (2018); Esterlis et al. (2021); Chubukov et al. (2004), in which gapless fermionic excitations are present in vicinity of the transition point, are ubiquitous to strongly correlated quantum materials such as heavy-fermion systems Paschen and Si (2021); Stewart (1984); Si et al. (2001); Custers et al. (2003); Senthil et al. (2004), cuprate and Fe-based superconductors Keimer et al. (2015); Lee et al. (2006); Sachdev (2003); Chen et al. (2014); Lee (2018b); Fernandes et al. (2022), and twisted bilayer graphene Cao et al. (2018a, b); Jaoui et al. (2022); Andrei and MacDonald (2020). In the aspects of equilibrium nature, many intriguing phenomena are unveiled arising from the interplay between critical fluctuations of order parameter and gapless fermions, including strange metallic transport Prochaska et al. (2020); Patel et al. (2023); Shen et al. (2020); Varma (2020); Phillips et al. (2022), high-$T_{c}$ superconductivity Scalapino (2012); Lederer et al. (2015); Li et al. (2016a); Berg et al. (2012); Marel et al. (2003) and intertwined ordering Kivelson et al. (2003); Fradkin et al. (2015); Schattner et al. (2016); Wang et al. (2018); Li et al. (2017a); Gerlach et al. (2017); Varma (1999). In contrast, the nonequilibrium properties of quantum driven dynamics in fermionic QCP are largely unaddressed.
The QCP occuring in Dirac semimetal (DSM), dubbed as Dirac QCP, represents a typical class of fermionic QCP. The studies of Dirac QCP stem from the research in modern high-energy physics, such as chiral symmetry in QCD and mass generation via spontaneously symmetry breaking Gross and Neveu (1974); Roy et al. (2016); Gracey et al. (2016); Gehring et al. (2015); Poland et al. (2019); Gazit et al. (2017); You et al. (2018). From the perspectives of statistical mechanics and condensed matter physics, Dirac quantum criticality also attracts enormous attentions, particularly after the experimental realization of two-dimensional Dirac fermions in graphene and various topological insulators or semimetals. The presence of Dirac fermions is theoretically revealed to tremendously enrich quantum critical properties, rendering novel universality classes of quantum phase transition without classical counterpart Ladovrechis et al. (2023); Janssen and Herbut (2014); Parisen Toldin et al. (2015); Knorr (2016); Ihrig et al. (2018); Boyack et al. (2021); Lang and Läuchli (2019); Liu et al. (2023); Li et al. (2018); Zhou et al. (2016); Tabatabaei et al. (2022); Otsuka et al. (2018); Xu and Grover (2021); Chandrasekharan and Li (2013); Wang and Meng (2023); Yu et al. (2024); Xu et al. (2023); Zhu et al. (2022); Sorella and Tosatti (1992); Herbut (2006); Herbut et al. (2009); Assaad and Herbut (2013); Otsuka et al. (2016); Mihaila et al. (2017); Zerf et al. (2017); Grover et al. (2014); Seifert et al. (2020); Ray and Janssen (2021); Corboz et al. (2018); Rader and Läuchli (2018); Da Liao et al. (2022); Wang et al. (2014); Li et al. (2015a); Liu et al. (2021, 2022); Chen et al. (2019); Zhang et al. (2019); Roy and Das Sarma (2016); Han et al. (2018); Herbut (2023); Otsuka et al. (2020); Li et al. (2017b, 2020); Torres et al. (2018); Classen et al. (2017); Yin and Zuo (2020). Consequently, fathoming the non-equilibrium driven dynamics in Dirac QCP has overarching meaning in basic theory, as well as immediate applications in the context of detecting and exploring fermionic QCP in experimental platform including realistic materials and quantum devices Castro Neto et al. (2009); Hasan and Kane (2010); Qi and Zhang (2011). Particularly, whether the original KZM is still applicable in the presence of gapless fermions remains largely elusive, constituting a fundamental question that bears profound impact on the theory of phase transition and nonequilibrium physics.
However, directly attacking the real-time dynamics in two or higher spatial dimension is largely hindered by the lack of reliable theoretical or numerical methods. Specifically, quantum Monte Carlo (QMC) fails as a result of the notorious sign problem Assaad and Evertz (2008); Troyer and Wiese (2005); Li and Yao (2019), while the tensor-network method still needs tremendous improvements despite remarkable progress in recent years Schmitt et al. (2022); Dziarmaga and Mazur (2023). Fortunately, scaling analysis demonstrates that both real- and imaginary-time driven dynamics share the same scaling form De Grandi et al. (2011); Grandi et al. (2013); Liu et al. (2013). This inference has been verified in various systems Liu et al. (2013); Shu et al. (2023); Schmitt et al. (2022), and particularly, substantiated by the remarkable consistency between the imaginary-time scaling and the annealing experiments in quantum devices Liu et al. (2015); King et al. (2023), bridging the gap between the QMC imaginary-time simulations without sign problem and the real-time dynamics.
In this work, for the first time we investigate the driven critical dynamics of two representative Dirac quantum critical points, belonging to chiral Heisenberg and chiral Ising universality classes respectively, via the determinant QMC method Assaad and Evertz (2008). By linearly varying the interaction strength along the imaginary-time direction to cross the QCP from both the DSM and Mott insulator phases, we uncover that for large driving rate the quantum critical dynamics is governed by the driving rate. The numerical results confirm that the whole driven process satisfies the scaling form of FTS. Hence, our study significantly generalizes the conventional KZM theory. Since DSM phase harbors gapless excitations, our results unambiguously show that the prerequisite of KZM theory that the initial state is fully gapped can be relaxed to the systems accommodating gapless Dirac fermions. In addition, our numerical simulation achieves the critical exponents of phase transition, whose values obtained in previous studies on equilibrium properties are still under debate. Our approach of QMC simulation on driven critical dynamics offers a novel theoretical tool to unravel properties of Dirac QCP. Through the innovative generalization of quantum driven critical dynamics to Dirac-fermion systems, our study not only leads to a great leap to the fundamental theory of KZM and FTS, but also contributes to a feasible approach to investigate the fermionic quantum critical phenomena in realistic platform such as quantum materials and devices.
Dynamics in chiral Heisenberg criticality— A typical model hosting QCP belonging to chiral Heisenberg universality class is the Hubbard model on the half-filled honeycomb lattice with the Hamiltonian Sorella and Tosatti (1992); Herbut (2006); Herbut et al. (2009); Assaad and Herbut (2013); Otsuka et al. (2016)
$H=-t\sum_{\langle ij\rangle,\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}+U\sum_{i}\left({n_{i\uparrow}-\frac{1}{2}}\right)\left({n_{i\downarrow}-\frac{1}{2}}\right),$ (1)
[…]
Figure 1: Sketch of the phase diagram in Dirac systems and the protocol for driven dynamics with different initial states. The correlation time scales for both boson (yellow curve) and fermion (violet curve) are finite in one phase (solid) but divergent (dotted) in the other phase. Accordingly, the prerequisite of the original KZM that a gapped initial state should exist to protect an initial adiabatic stage, in which the transition time is larger than (dashed line) than the correlation time, breaks down.
For the driven dynamics across the critical point, based on an adiabatic-impulse-adiabatic scenario, the original KZM often requires a gapped initial state to maintain an adiabatic initial evolution Dziarmaga (2010); Polkovnikov et al. (2011); Polkovnikov and Gritsev (2008), since the frozen time, which is a central notion in the KZM, lives at the edge of the adiabatic stage. After that, the KZM predicts an impulse regime wherein the system ceases evolving due to the critical slowing down Zurek (1985); Dziarmaga (2010); Polkovnikov et al. (2011). This oversimplified assumption was further improved by the FTS theory, which demonstrated that a driven-induced time scale controls the dynamics in the impulse region and the whole evolution process can be described by the FTS forms Gong et al. (2010); Feng et al. (2016); Huang et al. (2014). Here we begin to explore whether the FTS forms are still applicable in Dirac systems with gapless initial states.
Figure 2: Driven dynamics from the DSM phase of model (1). (a) Log-log plots of $m^{2}$ versus $R$ for different $L$ driven to $U_{c}=3.85$ before (a1) and after (a2) rescaling. Inset in (a1) shows $m^{2}\propto L^{-2}$ for $R=0.15$ (dash-dotted line). For large $R$, power fitting for $L=24$ (black solid line) shows $m^{2}\propto R^{-0.26}$ with the exponent close to $(2\beta-d\nu)/\nu r=-0.257$ (dash line) from Ref. Otsuka et al. (2016). (b) Curves of $m^{2}$ versus $g$ for fixed $RL^{r}=5.41$ and different $L$ before (b1) and after (b2) rescaling.
First we study the driven dynamics by varying $U$ with imaginary time $\tau$ as $U=U_{0}+R\tau$ starting from the DSM initial state with interaction strength $U_{0}$, as illustrated in Fig. 1. We denote the distance to the critical point as $g$ (Here $g=U-U_{c}$). When $g=0$, from Fig. 2 (a1), we find that for large $R$, $m^{2}\propto R^{(2\beta-d\nu)/\nu r}$ in which $\beta$ and $\nu$ are the exponents for order parameter and correlation length respectively Otsuka et al. (2016), and $r=z+1/\nu$ with $z$ the dynamic exponent. And the exponent on $R$ is almost independent of $L$. In addition, for fixed $R$, we find that $m^{2}\propto L^{-2}$. Combination of these results yields the scaling relation $m^{2}\propto L^{-d}R^{(2\beta-d\nu)/\nu r}$. In contrast, when $R$ is small, Fig. 2 (a1) shows that $m^{2}$ tends to saturate and the usual finite-size scaling $m^{2}\propto L^{-2\beta/\nu}$ is restored. To reconcile these rescaling relations, the scaling form must satisfy
$m^{2}(R,L,g)=L^{-d}R^{(2\beta-d\nu)/\nu r}\mathcal{F}(RL^{r},gL^{1/\nu}),$ (2)
in which $\mathcal{F}$ is a non-singular scaling function and therein dimensionless quantity $gL^{1/\nu}$ is also included to take account of the off-critical-point effects. Eq. (2) is consistent with the FTS in conventional bosonic QCP Huang et al. (2014); Liu et al. (2014).
To confirm Eq. (2), we find that the Eq. (2) yields the scaling form $m^{2}=L^{-2\beta/\nu}\mathcal{F}(RL^{r},0)$ with fixing $g=0$. Here, we rescale $m^{2}$ and $R$ as $m^{2}L^{2\beta/\nu}$ and $RL^{r}$ respectively, and reveal that the rescaled curves collapse well into a single curve, as shown in Fig. 2 (a2). The results of data collapse not only confirm Eq. (2) at $g=0$, but also verify the values of critical point and critical exponents.
Figure 3: Driven dynamics from the AFM phase of model (1). (a) Log-log plots of $m^{2}$ versus $R$ for different $L$ driven to $U_{c}=3.85$ before (a1) and after (a2) rescaling. For large $R$, power fitting (black solid line) shows $m^{2}\propto R^{0.71}$ with the exponent close to $0.7522\beta/\nu r=$ 0.752 (dash line) from Ref. Otsuka et al. (2016). (b) Curves of $m^{2}$ versus $g$ for fixed $RL^{r}=41.5$ and different $L$ before (b1) and after (b2) rescaling.
In addition, to unravel the scaling properties in the whole driven process, we calculate the dependence of $m^{2}$ on $g$ for an arbitrary fixed $RL^{r}$ and present the results in Fig. 2 (b1). After rescaling $m^{2}$ and $g$ by $L^{2\beta/\nu}$ and $L^{1/\nu}$ respectively, the curves with various $L$ collapse into each other, as displayed in Fig. 2 (b2), confirming that the universal scaling behavior of physical observable in the whole driven process is described by Eq. (2).
The reason for the appearance of the scaling relation $m^{2}\propto L^{-d}R^{(2\beta-d\nu)/\nu r}$ is that for large $R$, driven induced length scale $\xi_{R}\sim R^{-1/r}$ is smaller than $L$. Thus, the definition of $m^{2}$ indicates that $m^{2}\propto L^{-d}$ owing to the central limit theorem. Meanwhile, the rest part of the dimension of $m^{2}$ should be borne by $R$, giving rise to the leading term of Eq. (2). In this case, $\mathcal{F}(x,0)$ tends to a constant. In contrast, for small $R$, $\xi_{R}>L$ and the conventional finite-size scaling at equilibrium $m^{2}\propto L^{-2\beta/\nu}$ is recovered, as confirmed in Fig. 2 (a2), and $\mathcal{F}(x,0)$ obeys the scaling form $\mathcal{F}(x,0)\sim x^{d/r-2\beta/\nu r}$.
Figure 4: Driven dynamics from the DSM phase of model (4). (a) Log-log plots of $m^{2}$ versus $R$ for different $L$ driven to $V_{c}=1.355$ before (a1) and after (a2) rescaling. Inset in (a1) shows $m^{2}\propto L^{-2}$ for $R=1$ (dash-dotted line). For large $R$, power fitting for $L=24$ (black solid line) shows $m^{2}\propto R^{-0.32}$ with the exponent close to $(2\beta-d\nu)/\nu r=-0.31$ (dash line) from Ref. Hesselmann and Wessel (2016). (b) Curves of $m^{2}$ versus $g$ for fixed $RL^{r}=63.34$ and different $L$ before (b1) and after (b2) rescaling.
Next, we turn to explore the driven dynamics starting from the Mott insulator initial state. This state has the AFM order with transverse gapless modes. For large $R$, Fig. 3 (a1) shows that $m^{2}$ obeys $m^{2}\propto R^{2\beta/\nu r}$ and is nearly independent of $L$. Combining this scaling relation with the usual finite-size scaling $m^{2}\propto L^{-2\beta/\nu}$ which is restored for small $R$, the scaling form should obey
$m^{2}(R,L,g)=R^{2\beta/\nu r}\mathcal{G}(RL^{r},gL^{1/\nu}),$ (3)
where $\mathcal{G}$ is the non-singular scaling function and $g$ is also included therein. Eq. (3) is also accordant with the conventional FTS with ordered initial state Gong et al. (2010); Feng et al. (2016); Huang et al. (2014)
We rescale curves of $m^{2}$ versus $R$ for various $L$ at $g=0$ according to the scaling function $m^{2}(R,L,g)=L^{-2\beta/\nu}\mathcal{G}(RL^{r},0)$ and find that the rescaled curves collapse well, which confirms Eq. (3) at $g=0$. Furthermore, Fig. 3 (b1) shows the curves of $m^{2}$ versus $g$ for an arbitrary fixed $RL^{r}$. The rescaled results $(gL^{1/\nu},m^{2}L^{2\beta/\nu})$ for various $L$ collapse into a single smooth curve, as displayed in Fig. 3 (b2), confirming that the driven process from AFM initial state is described by Eq. (3).
The appearance of $m^{2}\propto R^{2\beta/\nu r}$ reflects the fact that when $\xi_{R}<L$ the initial ordered magnetization domain is maintained. In this case, $\mathcal{G}(x,0)$ tends to a constant. In contrast, for small $R$ with $\xi_{R}>L$, the usual finite size scaling is recovered, indicating $\mathcal{G}(x,0)\sim x^{-2\beta/\nu r}$ as $x\rightarrow 0$.
Figure 5: Driven dynamics from the CDW phase of model (4). (a) Log-log plots of $m^{2}$ versus $R$ for different $L$ driven to $V_{c}=1.355$ before (a1) and after (a2) rescaling. For large $R$, power fitting (black solid line) shows $m^{2}\propto R^{0.50}$ with the exponent close to $0.542\beta/\nu r=$ 0.54 from Ref. Hesselmann and Wessel (2016) (dash line). (b) Curves of $m^{2}$ versus $g$ for fixed $RL^{r}=87.97$ and different $L$ before (b1) and after (b2) rescaling.
Dynamics in chiral Ising criticality— To further verify the FTS in Dirac systems, we also explore the driven dynamics of Dirac QCP belonging to chiral Ising universality class, which is realized in the interacting spinless fermion model on the half-filled honeycomb lattice with the Hamiltonian Wang et al. (2014); Li et al. (2015a); Wang et al. (2016); Hesselmann and Wessel (2016); Hesselmann et al. (2018):
$H=-t\sum_{\langle ij\rangle}c_{i}^{\dagger}c_{j}+V\sum_{\langle ij\rangle}\left({n_{i}-\frac{1}{2}}\right)\left({n_{j}-\frac{1}{2}}\right),$ (4)
where $V$ measures the nearest-neighbor interaction. The model is amendable to sign-problem-free QMC simulationLi et al. (2015b, 2016b); Wei et al. (2016); Wang et al. (2015). It is shown that the ground state undergoes a continuous quantum phase transition at $V_{c}\approx 1.355t$ from the DSM phase to the charge-density-wave (CDW) insulating phase, characterized by the order parameter $m^{2}\equiv\sum_{i,j}\eta_{i}\eta_{j}\langle{(n_{i}-1/2)(n_{j}-1/2)}\rangle/L^{2d}$ with $\eta_{i}=\pm 1$ for $i$in $A(B)$ sublattice Wang et al. (2014); Li et al. (2015a); Hesselmann and Wessel (2016). At $V=V_{c}$, both fermion and boson degrees of freedom are gapless, similar to model (1). However, a difference is that in the CDW phase, the bosonic fluctuation is fully gapped owing to the discrete symmetry breaking.
From the DSM initial state, Fig. 4 (a1) shows that fixed at $V=V_{c}$, namely $g=0$, $m^{2}\propto L^{-d}R^{(2\beta-d\nu)/\nu r}$ for large $R$, whereas $m^{2}\propto L^{-2\beta/\nu}$ for small $R$, similar to the results in model (1). In addition, we find the rescaled curves of $m^{2}$ versus $R$ and $g$ collapse well, as shown in Fig. 4 (a2) and (b2), confirming that Eq. (2) gives a universal description on the driven critical dynamics from the DSM initial state.
From the CDW initial state, Fig. 5 (a1) shows that at $g=0$, $m^{2}\propto R^{2\beta/\nu r}$ for large $R$, whereas $m^{2}\propto L^{-2\beta/\nu}$ for small $R$, similar to the case of chiral Heisenberg universality class. Moreover, Eq. (3) is verified by the data collapse of the curves $(RL^{r},m^{2}L^{-2\beta/\nu})$ at fixed $g=0$ and $(gL^{1/\nu},m^{2}L^{-2\beta/\nu})$ for an arbitrary $RL^{r}$, as shown in Fig. 5 (a2) and (b2). Consequently, Eq. (3) provides a universal description on the driven dynamics from ordered initial states, regardless of whether or not gapless bosonic modes exist.
Discussion— Some remarks are discussed as follows. (a) A crucial prerequisite in the original KZM is the existence of initial adiabatic stage in which the excitations possess finite gap Dziarmaga (2010); Polkovnikov et al. (2011); Polkovnikov and Gritsev (2008). In contrast, the FTS of driven critical dynamics still holds in Dirac systems, indicating that this requirement is not indispensable, since in both quantum phases separated by the QCP the gapless excitations are present: the DSM phase features gapless fermions, whereas bosonic gapless excitation exists in the AFM Mott insulator phase where continuous symmetry is spontaneously broken.
To understand the underlying reasons of the novel numerical results, we notice that the DSM phase has the vanishing density of states at charge neutral point. Moreover, the Pauli principle of fermions further restricts the excitations in the vicinity of Dirac points. Accordingly, the excitations under slow external driving are strongly suppressed in DSM phase. Hence, it is speculated that the theories of KZM and FTS in driven dynamics are generally obeyed in Dirac QCP with the initial state of DSM. Besides, for driven dynamics from the AFM initial state, since we focus on the dynamic scaling of the order parameter magnitude, the influence of gapless angular excitation is tiny since its characteristic time scale is much longer than that of longitudinal mode Weinberg and Sandvik (2017), similar to the case in deconfined quantum criticality Shu et al. (2023).
(b) Furthermore, our work provides an effective approach to assess fermionic critical properties from the perspective of nonequilibirum driven dynamics. Large discrepancy on critical exponents of chiral Heisenberg universality class still exists even under extensive studies Assaad and Herbut (2013); Otsuka et al. (2016). Here, by means of FTS analysis we reveal that the critical exponents for large $R$ in Figs. 2 and 3 are consistent with the ones from Ref. Otsuka et al. (2016), in which results of larger system size and careful finite-size corrections are included. For the finite-size scaling of equilibrium QCP, the critical exponents are extracted by the scaling analysis with varying interaction strength and system size. In the framework of driven dynamics, the driven rate $R$ constitutes a new tuning parameter, enabling the achievement of critical exponents with fixing $g=0$ or $L=L_{\rm max}$ according to Eqs. (2) and (3), which offers a possible way to mitigating sub-leading corrections arising from finite $g$ or small system size. Consequently, the reliable values of critical exponents are accessible with relatively small system size.
Summary— In summary, we investigate the driven dynamics of QCP in two representative interacting Dirac-fermion systems, belonging to chiral Heisenberg and chiral Ising universality classes respectively, through sign-problem-free QMC simulation. Driving the system from both DSM and Mott insulator phase as the initial state, we discover varieties of interesting nonequilibrium scaling behaviors. Furthermore, we confirm that these scaling behaviors can be unified by the full scaling form of the FTS theory. Through this work we not only successfully generalize the KZM and FTS to critical systems with joint fluctuations of gapless fermions and bosons, but also extend the application conditions of the celebrated theory of KZM. Additionally, we demonstrate that the nonequilibrium scaling form is capable of determining the critical exponents in Dirac QCP, providing an effective method to deciphering quantum critical properties in terms of driven dynamics. Our work possesses many intriguing generalizations, for example, it is quite instructive to consider other kinds of driven protocols, such as driving from the critical point and from the thermal equilibrium state Chandran et al. (2012); Deng et al. (2009); De Grandi et al. (2010); Huang et al. (2016); Yin et al. (2016). From the experimental perspectives, the Moiré systems of graphene and heterostructures offer tunable platforms to observe the fermionic QCP and the associated driven critical dynamicsKennes et al. (2021); Li et al. (2022). Moreover, as aforementioned, the Kibble-Zurek driven dynamics is experimentally observed in cold-atom systemsClark et al. (2016); Keesling et al. (2019). Owing to the recent developments of cold atoms based quantum simulator of fermions Jördens et al. (2008); Mazurenko et al. (2017); Venu et al. (2023), it is promising to detect driven dynamics in Dirac QCP and verify the generalized KZM and FTS as discussed in our study in these platforms.
Acknowledgments— We would like to thank A. W. Sandvik for helpful discussions. Z. Zeng, Y. K. Yu, Z. X. Li and S. Yin are supported by the National Natural Science Foundation of China (Grants No. 12075324 and No. 12222515). Z. X. Li is supported by the NSFC under Grant No. 12347107. S. Yin is also supported by the Science and Technology Projects in Guangdong Province (Grants No. 211193863020). The calculations reported were partly performed on resources provided by the Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices, No. 2022B1212010008.
Supplementary Materials for
Finite-time Scaling beyond the Kibble-Zurek Prerequisite: Driven Critical Dynamics in Strongly Interacting Dirac Systems
I IMAGINARY-TIME DRIVEN DYNAMICS
The driven dynamics is realized by varying $U$ with imaginary time $\tau$ as $U=U_{0}+R\tau$, starting from the ground state at $U_{0}$. The wave function $|\psi(\tau)\rangle$ obeys the imaginary-time Schrödinger equation De Grandi et al. (2011); Grandi et al. (2013); Liu et al. (2013)
$-\frac{\partial}{\partial\tau}|\psi(\tau)\rangle=H(\tau)|\psi(\tau)\rangle.$ (S1)
[…]
II Determinant quantum Monte Carlo
We employ the large-scale determinant QMC method Assaad and Evertz (2008) to investigate the imaginary-time driven dynamics of our models. Specifically, we start with a trial wave function |$\ket{\psi_{T}}$ , and project the system onto the desired initial state |$\ket{\psi_{0}}$ with a sufficiently long projection time $\tau_{0}$. Then we vary the system parameter linearly with imaginary time variable $\tau$ and observe the scaling behavior of observables at final state, at which the expectation value of observables is given by
$\braket{O(\tau)}=\frac{\bra{\psi_{0}}U(2\tau,\tau)~{}O~{}U(\tau,0)\ket{\psi_{0}}}{\bra{\psi_{0}}U(2\tau,0)\ket{\psi_{0}}}=\frac{\bra{\psi_{T}}\mathrm{e}^{-\tau_{0}H_{\mathrm{initial}}}U(2\tau,\tau)~{}O~{}U(\tau,0)\mathrm{e}^{-\tau_{0}H_{\mathrm{initial}}}\ket{\psi_{T}}}{\bra{\psi_{T}}\mathrm{e}^{-\tau_{0}H_{\mathrm{initial}}}U(2\tau,0)\mathrm{e}^{-\tau_{0}H_{\mathrm{initial}}}\ket{\psi_{T}}}.$ (S2)
In numerical calculations, we use Trotter decomposition to discretize imaginary-time propagator into ${M=\Theta/\Delta\tau}$ (${M}$ is integer and $\Theta$ is imaginary time) time slices with
$\mathrm{e}^{-\Theta H}=\prod^{M}_{m=1}\left[\mathrm{e}^{-\Delta\tau H_{t}}\mathrm{e}^{-\Delta\tau H_{U}}+\mathcal{O}\left(\Delta\tau^{2}\right)\right],$ (S3)
where $H_{t}$ and $H_{U}$ are the hopping term and the interaction term respectively in the Hamiltonian. We choose small enough $\Delta\tau/t\leq 0.05$. To decouple two-body fermion-fermion coupling form of $\mathrm{e}^{\Delta\tau H_{U}}$, we use discrete Hubbard-Stratonovich transformation
$\mathrm{e}^{-\frac{\Delta\tau U}{2}\left(n_{i\uparrow}+n_{i\downarrow}-1\right)^{2}}=\sum_{l=\pm 1,\pm 2}\gamma(l)\mathrm{e}^{\mathrm{i}\sqrt{\frac{\Delta\tau U}{2}}\eta(l)\left(n_{i\uparrow}+n_{i\downarrow}-1\right)},$ (S4)
$\mathrm{e}^{-\frac{\Delta\tau V}{2}\left(n_{i}+n_{j}-1\right)^{2}}=\sum_{l=\pm 1,\pm 2}\gamma(l)\mathrm{e}^{\mathrm{i}\sqrt{\frac{\Delta\tau V}{2}}\eta(l)\left(n_{i}+n_{j}-1\right)},$ (S5)
to obtain one-body fermion-auxiliary field coupling. Here, we introduce four-component space-time local auxiliary fields $\gamma(\pm 1)=1+\sqrt{6}/3$, $\gamma(\pm 2)=1-\sqrt{6}/3$, $\eta(\pm 1)=\pm\sqrt{2\left(3-\sqrt{6}\right)}$, $\eta(\pm 2)=\pm\sqrt{2\left(3+\sqrt{6}\right)}$, and use DQMC for importance sampling over these space-time configurations. Next, we elaborate on how DQMC numerically calculates the sampling weight.
Taking Hubbard interaction as an example, for each imaginary time and each position of the interaction, we employ Hubbard-Stratonovich transformation as in Eq. (S4). This means that we introduce an auxiliary field in $d+1$ dimensions. As a result, the imaginary-time propagator can be fully expressed using single-particle operators. This allows us to represent it in the following quadratic form of fermion operators:
$\mathrm{e}^{-\Theta H}\equiv\sum_{\mathrm{c}}\mathrm{e}^{-\Theta H_{\mathrm{c}}}=\sum_{\mathrm{c}}A_{\mathrm{c}}\prod^{M}_{m=1}~{}\mathrm{e}^{\vec{c}^{\dagger}T\vec{c}}~{}\mathrm{e}^{\vec{c}^{\dagger}V_{\mathrm{c}(m)}\vec{c}},$ (S6)
where $\sum_{\mathrm{c}}$ denotes the summation over all space-time configurations of the auxiliary field. Considering that each local component of the auxiliary field has 4 possible values, the summation comprises up to $4^{MN}$ terms, where $N$ represents the number of spatial degrees of freedom. $H_{c}$ denotes the decoupled configuration Hamiltonian, while $T$ and $V_{\mathrm{c}(m)}$ are the resulting quadratic coefficient matrices from the rearrangement, and $A_{\mathrm{c}}$ is the coefficient. Both $V_{\mathrm{c}(m)}$ and $A_{\mathrm{c}}$ depend on the auxiliary field configuration. The complete form of the evolution operator has been presented above. Next, we consider expressing the trial wave function. We use a DSM inital state for trial wave function, which is a direct product state, and numerically it can be written as the following Slater determinant:
|$\ket{\psi_{T}}=\bigotimes^{N_{\mathrm{e}}}_{n_{\mathrm{e}}=1}\left[\left(\sum_{x}c_{x}^{\dagger}P_{x,n_{\mathrm{e}}}\right)\ket{0}\right]=\bigotimes^{N_{\mathrm{e}}}_{n_{\mathrm{e}}=1}\left[\left(\vec{c}^{\dagger}P\right)_{n_{\mathrm{e}}}\ket{0}\right],$ (S7)
where $N_{\mathrm{e}}$ denotes the number of electrons. This implies that the trial wave function is a direct product of $N_{\mathrm{e}}$ fermion single-particle wave functions. The index $x$ denotes the degree of freedom of the electron, including spatial degrees of freedom, spin degrees of freedom, etc. The matrix element $P_{x,n_{\mathrm{e}}}$ represents the probability amplitude of the $n_{\mathrm{e}}$th electron on the $x$th degree of freedom. Note that the imaginary-time propagator $\mathrm{e}^{-\Theta H_{\mathrm{c}}}$ is essentially the Boltzmann weight factor of the auxiliary field configuration in statistical mechanics. According to Eqs. (S6) and (S7), the partition function of the auxiliary field configuration can be expressed as:
$Z=\sum_{\mathrm{c}}\bra{\psi_{T}}\mathrm{e}^{-\Theta H_{\mathrm{c}}}\ket{\psi_{T}}=\sum_{\mathrm{c}}A_{\mathrm{c}}~{}\mathrm{det}\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right].$ (S8)
Here, we use $B_{\mathrm{c}}$ to represent the exponential of the quadratic coefficient matrix:
$B_{\mathrm{c}}(\tau_{2},\tau_{1})\equiv\prod^{\tau_{2}/\Delta\tau}_{m=\tau_{1}/\Delta\tau}~{}\mathrm{e}^{T}~{}\mathrm{e}^{V_{\mathrm{c}(m)}}.$ (S9)
The expression on the right side of Eq. (S8) has integrated out the fermion operators, replacing the Grassmann numbers and fermion statistics with a determinant representation which is computationally tractable. All matrix operations can be performed directly on a computer.
Ultimately, our Monte Carlo sampling is conducted over space-time configurations. Numerically, the weight of a space-time configuration is $A_{\mathrm{c}}~{}\mathrm{det}\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right]$. Following the classical Markov importance sampling method, we continuously make tentative flips to the local components of this $d+1$ dimensional auxiliary field. We then employ the Metropolis algorithm to calculate the probability of accepting these changes based on the ratio of configuration weights before and after the flip. Specifically, we need to compute the following weight ratio:
$R_{\mathrm{c^{\prime}c}}\equiv\frac{A_{\mathrm{c^{\prime}}}~{}\mathrm{det}\left[P^{\dagger}B_{\mathrm{c^{\prime}}}(\Theta,0)P\right]}{A_{\mathrm{c}}~{}\mathrm{det}\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right]},$ (S10)
where c represents the flipped configuration and c represents the original configuration. In fact, we do not need to compute the weights of the two configurations separately. This is because the flipping we perform is localized in space-time, so
$B_{\mathrm{c^{\prime}}}(\Theta,0)=B_{\mathrm{c}}(\Theta,\zeta)\left(\mathbf{1}+\Delta_{\mathrm{c^{\prime}c}}\right)B_{\mathrm{c}}(\zeta,0).$ (S11)
Here, $\Delta_{\mathrm{c^{\prime}c}}$ is a highly sparse matrix, where only the matrix elements corresponding to the degrees of freedom involved in the local auxiliary field flipping are non-zero. Thus, the ratio of the two determinants can be expressed as:
$\frac{\mathrm{det}\left[P^{\dagger}B_{\mathrm{c^{\prime}}}(\Theta,0)P\right]}{\mathrm{det}\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right]}=\mathrm{det}\left\{\mathbf{1}+\Delta_{\mathrm{c^{\prime}c}}B_{\mathrm{c}}(\zeta,0)P\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right]^{-1}P^{\dagger}B_{\mathrm{c}}(\Theta,\zeta)\right\}.$ (S12)
Due to the sparsity of $\Delta_{\mathrm{c^{\prime}c}}$, the determinant on the right side of the above equation only requires consideration of a few degrees of freedom involved in the flipping during computations.
In DQMC, to compute the physical observables, we only need to statistically analyze the configurational observable $\braket{O(\tau)}_{\mathrm{c}}$.
$\braket{O(\tau)}=\sum_{\mathrm{c}}\mathrm{Pr_{c}}\braket{O(\tau)}_{\mathrm{c}}+\mathcal{O}\left(\Delta\tau^{2}\right),$ (S13)
where $\mathrm{Pr_{c}}$ represents the configuration probability,
$\mathrm{Pr_{c}}=\frac{1}{Z}A_{\mathrm{c}}~{}\mathrm{det}\left[P^{\dagger}B_{\mathrm{c}}(2\tau+\tau_{0},0)P\right],$ (S14)
$\braket{O(\tau)}_{\mathrm{c}}=\frac{\bra{\psi_{0}}U_{\mathrm{c}}(2\tau,\tau)~{}O~{}U_{\mathrm{c}}(\tau,0)\ket{\psi_{0}}}{\bra{\psi_{0}}U_{\mathrm{c}}(2\tau,0)\ket{\psi_{0}}}=\frac{\bra{\psi_{T}}\mathrm{e}^{-\tau_{0}H_{\mathrm{initial},\mathrm{c}}}U_{\mathrm{c}}(2\tau,\tau)~{}O~{}U_{\mathrm{c}}(\tau,0)\mathrm{e}^{-\tau_{0}H_{\mathrm{initial},\mathrm{c}}}\ket{\psi_{T}}}{\bra{\psi_{T}}\mathrm{e}^{-\tau_{0}H_{\mathrm{initial},\mathrm{c}}}U_{\mathrm{c}}(2\tau,0)\mathrm{e}^{-\tau_{0}H_{\mathrm{initial},\mathrm{c}}}\ket{\psi_{T}}}.$ (S15)
Since we employ importance sampling, the sampling frequency is proportional to the configuration probability. Ultimately, when calculating the observable, we simply take the average over the sampled configurational observables. If the observable is a single-particle operator, meaning it can be expressed as a quadratic form of fermion operators, then one can integrate out the fermion degrees of freedom in a manner similar to Eq. (S8) and numerically compute using determinants. For observables of four-fermion operators or higher, we compute using the fermion equal-time Green’s function based on Wick’s theorem. After numerically integrating out the fermion degrees of freedom, the fermion equal-time Green’s function can be expressed using the following matrix element:
$\braket{c_{x_{1}}^{\dagger}c_{x_{2}}}_{\mathrm{c}}=\left\{B_{\mathrm{c}}\left(\frac{\Theta}{2},0\right)P\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right]^{-1}P^{\dagger}B_{\mathrm{c}}\left(\Theta,\frac{\Theta}{2}\right)\right\}_{x_{1},x_{2}}.$ (S16)
III RESULTS OF DIMENSIONLESS QUANTITY
In the main text, we only present dynamic results of the square of the order parameter $m^{2}$. Here, we supplement the results of a dimensionless quantity correlation ratio defined as Kaul (2015):
$R_{S}\equiv{1-S(\Delta\bm{q})}/{S(\bm{0})},$ (S17)
where $\Delta\bm{q}$ is the minimum lattice momentum and $S({\bm{q}})$ is the structure factor.
Figure S1: Log-log plots of $R_{S}$ versus $R$ for different $L$ driven to $U_{c}=3.85$, for model (1). (a1) displays the results driven from the DSM phase. (a2) shows rescaling is applied to the horizontal axis of (a1). (b1) displays the results driven from the AFM phase. (b2) shows rescaling is applied to the horizontal axis of (b1).
Figure S2: Log-log plots of $R_{S}$ versus $R$ for different $L$ driven to $V_{c}=1.355$, for model (4). (a1) displays the results driven from the DSM phase. (a2) shows rescaling is applied to the horizontal axis of (a1). (b1) displays the results driven from the CDW phase. (b2) shows rescaling is applied to the horizontal axis of (b1).
For model (1), $S({\bm{q}})$ is the antiferromagnetic structure factor:
$S({\bm{q}})_{\rm AFM}=\frac{1}{L^{2d}}\sum_{i,j}\mathrm{e}^{\mathrm{i}\bm{q}\cdot({\bm{r}_{i}-\bm{r}_{j}})}\langle{S_{i}^{z}S_{j}^{z}}\rangle,$ (S18)
where $S_{i}^{z}$ is the staggered magnetization operator defined as $S_{i}^{z}\equiv\vec{c}^{\dagger}_{i,A}\sigma^{z}\vec{c}_{i,A}-\vec{c}^{\dagger}_{i,B}\sigma^{z}\vec{c}_{i,B}$ with $i$ characterizes cells and ${\vec{c}^{\dagger}\equiv({c_{\uparrow}^{\dagger},c_{\downarrow}^{\dagger}}})$.
For model (4), $S({\bm{q}})$ is the charge-density-wave structure factor:
$S({\bm{q}})_{\rm CDW}=\frac{1}{L^{2d}}\sum_{i,j}\mathrm{e}^{\mathrm{i}\bm{q}\cdot({\bm{r}_{i}-\bm{r}_{j}})}\langle{n_{i}n_{j}}\rangle,$ (S19)
where $n_{i}\equiv n_{i,A}-n_{i,B}$ with $n_{i,A}$ being the electron number operator defined as $n_{i,A}\equiv c^{\dagger}_{i,A}c_{i,A}$.
According to the scaling analysis in the main text, as a dimensionless variable, $R_{S}$ in the driven process obeys the following dynamic scaling form:
$R_{S}({g,R,L})=f(RL^{r},gL^{1/\nu}),$ (S20)
which indicates that (S20) at $g=0$ can be converted to $R_{S}=f(RL^{r})$.
As shown in Fig. S1(a2)(b2) and Fig. S2(a2)(b2), we find that the rescaled curves all collapse well. These results not only confirm (S20) at $g=0$, but also support the conclusions we draw in the main text.